OpenStudy (anonymous):

Medals!!! The human brain weighs about 1 lb for each 100 lb of body weight. What is the approximate weight of the brain of a person weighing 85 lb to the nearest ounce?

OpenStudy (anonymous):

let x be the weight of the brain \[\frac{ 1 }{ 100 }=\frac{ x }{ 85 }\] 100x=85*1 x=85/100 x=0.85lb

OpenStudy (anonymous):

about 14 ounces

OpenStudy (anonymous):

You guys both gave really good answers... :/ How about I become a fan of one of you, and give a medal to the other?

OpenStudy (anonymous):

Thank you both very much!

OpenStudy (anonymous):

your wlkome

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

thnx Abdulhameed